there are two coins. One is fair and the other one has a 5/8 probability to heads. A coin is chosen at random and tossed twice. Heads shows twice. What is the probability the coin you chose is the biased one
Answer #1

P(heads tossed twice) =P(fair coin and 2 heads)+P(biased coin and 2 heads)

=(1/2)*(1/2)*(1/2)+(1/2)*(5/8)*(5/8)= 0.3203

P(Biased coin given two heads) =P(biased coin and 2 heads)/P(2 heads)

=(1/2)*(5/8)*(5/8)/0.3203 =25/41=0.609756